OpenStudy (anonymous):

can someone help me find the absolute and local extrema in this equation (3x-4)/(x^2+1)

OpenStudy (amistre64):

sure

OpenStudy (amistre64):

(x^2+1)(3) - (3x-4)(2x) = 0

OpenStudy (anonymous):

on the intervals [-2,2]

OpenStudy (amistre64):

3x^2 +3 -6x^2 +8x = 0 -3x^2 +3 +8x = 0 0 = 3x^2 -8x -3

OpenStudy (amistre64):

(x-3) (x+ 1/3) = x = x and x = -1/3 are the max and min

OpenStudy (amistre64):

x = 3 and x = -1/3......

OpenStudy (amistre64):

f(-2) = ? f(-1/3) = ? f(2) = ?

OpenStudy (anonymous):

so what do u do with the denominator

OpenStudy (amistre64):

nothing; the top directs the =0 part; not the bottom

OpenStudy (amistre64):

put these x values into the original equationand determine which is higher

OpenStudy (anonymous):

awesome thanks a lot man the denominator was throwing me off

OpenStudy (amistre64):

there is no vertical asypmtote; so you aint gotta worry about bad values :)

myininaya (myininaya):

also x^2+1=0 has no real solution so this does not give you a critcal number anyways

myininaya (myininaya):

right what amistre said no vertical asym

OpenStudy (anonymous):

awesome thanks a lot guys